# QECO: A QoE-Oriented Computation Offloading Algorithm based on Deep Reinforcement Learning for Mobile Edge Computing

### v2

@@ II-B1 Local Execution
$l_{i}^{\text{C}}(t)=\min\Big{\{}t+\delta_{i}^{\text{C}}(t)+\lceil D_{i}^{\text{C}}(t)\rceil-1,t+\Delta_{i}(t)-1\Big{\}}.$ (6)
The task $z_{i}(t)$ will be immediately dropped if its processing is not completed by the end of the time slot $t+\Delta_{i}(t)-1$. In addition, we introduce $D_{i}^{\text{C}}(t)$ as the number of time slots required to complete the processing of task $z_{i}(t)$ on MD $i\in\mathcal{I}$. It is given by:
@@
II-B2 Edge Execution
We model the edge execution by the queues associated with MDs deployed at ENs. If computation task $z_{i}(t^{\prime})$ is dispatched to EN $j$ in time $t^{\prime}<t$, we let $z_{i,j}^{\text{E}}(t)$ and $\lambda_{i,j}^{\text{E}}(t)$ (in bits) denote the unique index of the task and the size of the task in the $i^{\text{th}}$ queue at EN $j$. We define $\eta_{i,j}^{\text{E}}(t)$ (in bits) as the length of this queue at the end of time slot $t\in\mathcal{T}$. We refer to a queue as an active queue in a certain time slot if it is not empty. That being said, if at least one task is already in the queue from previous time slots or there is a task arriving at the queue, that queue is active. We define $\mathcal{B}_{j}(t)$ to denote the set of active queues at EN $j$ in time slot $t$.
$\mathcal{B}_{j}(t)={\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\Big{\{}i{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\big{|}i\in\mathcal{I},\lambda_{i,j}^{\text{E}}(t)>0\,\,{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\text{or}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\,\,\eta_{i,j}^{\text{E}}(t-1)>0{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\Big{\}}.{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}$ (9)
@@ II-B2 Edge Execution
$\eta_{i,j}^{\text{E}}(t)=\left[\eta_{i,j}^{\text{E}}(t-1)+\lambda_{i,j}^{\text{E}}(t)-{f_{j}^{\text{E}}\tau\over\rho_{i}(t)B_{j}(t)}-\omega_{i,j}(t)\right]^{+}.$ (10)
@@
$\hat{l}_{i,j}^{\text{E}}(t)=\max\{t,\max\limits_{t^{{}^{\prime}}\in\{0,1,\ldots,t-1\}}l_{i,j}^{\text{E}}(t^{{}^{\prime}})+1\},$ (11)
@@
$\sum_{t^{{}^{\prime}}=\hat{l}_{i,j}^{\text{E}}(t)}^{l_{i,j}^{\text{E}}(t)}{f_{j}^{\text{E}}\tau\over\rho_{i}(t)B_{j}(t^{{}^{\prime}})}\mathbbm{1}(i\in\mathcal{B}_{j}(t^{{}^{\prime}}))\geq\lambda_{i,j}^{\text{E}}(t),$ (12)
$\sum_{t^{{}^{\prime}}=\hat{l}_{i,j}^{\text{E}}(t)}^{l_{i,j}^{\text{E}}(t)-1}{f_{j}^{\text{E}}\tau\over\rho_{i}(t)B_{j}(t^{{}^{\prime}})}\mathbbm{1}(i\in\mathcal{B}_{j}(t^{{}^{\prime}}))<\lambda_{i,j}^{\text{E}}(t),$ (13)
@@
$D_{i,j}^{\text{E}}(t)={\lambda_{i,j}^{\text{E}}(t)\rho_{i}(t)\over f_{j}^{\text{E}}\tau/B_{j}(t)}.$ (14)
@@
$E_{i,j}^{\text{E}}(t)={D_{i,j}^{\text{E}}(t)p_{j}^{\text{E}}\tau\over B_{j}(t)},$ (15)
where $p_{j}^{\text{E}}$ is a constant value which denotes the energy consumption of the EN $j$’s processor when operating at full capacity.
In addition to the energy consumed by EN $j$ for task processing, we also take into account the energy consumed by the MD $i$’s user interface in the standby state while waiting for task completion at the EN $j$. We define $E_{i,j}^{\text{I}}(t)$ as the energy consumption associated with the user interface of MD $i\in\mathcal{I}$, which is given by
$E_{i}^{\text{I}}(t)=D_{i,j}^{\text{E}}(t)p_{i}^{\text{I}}\tau,$ (16)
where $p_{i}^{\text{I}}$ is the standby energy consumption of MD $i\in\mathcal{I}$.
$E_{i}^{\text{O}}(t)=E_{i}^{\text{T}}(t)+\sum_{\mathcal{J}}E_{i,j}^{\text{E}}(t)+E_{i}^{\text{I}}(t).$ (17)
III Task Offloading problem Formulation
Based on the introduced system model, we present the computation task offloading problem in this section. Our primary goal is to enhance each MD’s QoE individually by taking the dynamic demands of MDs into account. To achieve this, we approach the optimization problem as an MDP, aiming to maximize the MD’s QoE by striking a balance among key QoE factors, including task completion, task delay, and energy consumption. To prioritize QoE factors, we utilize the MD’s battery level, which plays a crucial role in decision-making. Specifically, when an MD observes its state (e.g. task size, queue details, and battery level) and encounters a newly arrived task, it selects an appropriate action for that task. The selected action, based on the observed state, will result in enhanced QoE. Each MD strives to maximize its long-term QoE by optimizing the policy mapping from states to actions. In what follows, we first present the state space, action space, and QoE function, respectively. We then formulate the QoE maximization problem for each MD.
III-A State Space
A state in our MDP represents a conceptual space that comprehensively describes the state of an MD facing the environment. We represent the MD $i$’s state in time slot $t$ as vector $\boldsymbol{s}_{i}(t)$ that includes the newly arrived task size, the queues information, the MD’s battery level, and the workload history at the ENs. The MD observers this vector at the beginning of each time slot. The vector $\boldsymbol{s}_{i}(t)$ is defined as follows:
@@ III-A State Space
where vector $\boldsymbol{\eta}_{i}^{\text{E}}(t-1)=(\eta_{i,j}^{\text{E}}(t-1))_{j\in\mathcal{J}}$ represents the queues length of MD $i$ in ENs at the previous time slot and is computed by the MD according to (10). Let $\phi_{i}(t)$ denote the battery level of MD $i$ in time slot $t$. Considering the power modes of a real mobile device, $\phi_{i}(t)$ is derived from the discrete set $\Phi=\{\phi_{1},\phi_{2},\phi_{3}\}$, corresponding to ultra power-saving, power-saving, and performance modes, respectively.
In addition, to predict future EN workloads, we define the matrix $\mathcal{H}(t)$ as historical data, indicating the number of active queues for all ENs. This data is recorded over $T^{\text{s}}$ time slots, ranging from $t-T^{\text{s}}$ to $t-1$, in $T^{\text{s}}\times J$ matrix. For EN $j$ workload history at $i^{th}$ time slot from $T^{\text{s}}-t$, we define $h_{i,j}(t)$ as:
@@
III-B Action Space
The action space represents the agent’s behavior and the decisions. In this context, we define $\boldsymbol{a}_{i}(t)$ to denote the action taken by MD $i\in\mathcal{I}$ in time slot $t\in\mathcal{T}$. These actions involve two decisions, (a) Offloading decision to determine whether or not to offload the task, and (b) Offloading target to determine the EN to send the offloaded tasks. Thus, the action of MD $i$ in time slot $t$ can be concisely expressed as the following action tuple:
@@ III-B Action Space
where vector $\boldsymbol{y}_{i}(t)=(y_{i,j}(t))_{j\in\mathcal{J}}$ represents the selected EN for offloading this task. In Section IV-B, we will discuss about the size of this action space.
III-C QoE Function
The QoE function evaluates the influence of agent’s actions by taking several key performance factors into account. Given the selected action $\boldsymbol{a}_{i}(t)$ in the observed state $\boldsymbol{s}_{i}(t)$, we represent $\mathcal{D}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$ as the delay of task $z_{i}(t),$ which indicates the number of time slots from time slot $t$ to the time slot in which task $z_{i}(t)$ is processed. It is calculated by:
$\mathcal{D}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))=(1-x_{i}(t))\Big{(}l_{i}^{\text{C}}(t)-t+1\Big{)}+$
$\;\;\;\;x_{i}(t)\bigg{(}\sum\limits_{\mathcal{J}}\sum\limits_{t^{\prime}=t}^{T}\mathbbm{1}\big{(}z_{i,j}^{\text{E}}(t^{\prime})=z_{i}(t)\big{)}l_{i,j}^{\text{E}}(t^{\prime})-t+1\bigg{)},$ (21)
where $\mathcal{D}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))=0$ when task $z_{i}(t)$ is dropped. Correspondingly, we denote the energy consumption of task $z_{i}(t)$ when taking action $\boldsymbol{a}_{i}(t)$ in the observed state $\boldsymbol{s}_{i}(t)$ as $\mathcal{E}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$, which is:
$\mathcal{E}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))=(1-x_{i}(t))E_{i}^{\text{L}}(t)+$
$\hskip 51.21504ptx_{i}(t)\bigg{(}\sum\limits_{\mathcal{J}}\sum\limits_{t^{{}^{\prime}}=t}^{T}\mathbbm{1}\big{(}z_{i,j}^{\text{E}}(t^{\prime})=z_{i}(t)\big{)}E_{i}^{\text{O}}(t)\bigg{)}.$ (22)
Given the delay and energy consumtion of task $z_{i}(t)$, we also define $\mathcal{C}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$ that denotes the assosiate cost of task $z_{i}(t)$ given the action $\boldsymbol{a}_{i}(t)$ in the state $\boldsymbol{s}_{i}(t)$.
@@ III-C QoE Function
where $\phi_{i}(t)$ represents the MD $i$’s battery level. When the MD is operating in performance mode, the primary focus is on minimizing task delays, thus the delay contributes more to the cost. On the other hand, when the MD switches to ultra power-saving mode, the main attention is directed toward reducing power consumption.
Finally, we define $\boldsymbol{q}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$ as the QoE associated with task $z_{i}(t)$ given the selected action $\boldsymbol{a}_{i}(t)$ and the observed state $\boldsymbol{s}_{i}(t)$. The QoE function is defined as follows:
$\boldsymbol{q}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))=$
$\hskip 22.76219pt\begin{cases}\mathcal{R}-\mathcal{C}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))&\text{if task $z_{i}(t)$ is processed,}\\ -\hskip 2.27621pt\mathcal{E}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))&\text{if task $z_{i}(t)$ is dropped,}\end{cases}$ (24)

### v3

@@ III-B1 Local Execution
$l_{i}^{\text{C}}(t)=\min\bigg\{t+\delta_{i}^{\text{C}}(t)+\lceil D_{i}^{\text{C}}(t)\rceil-1,t+\Delta_{i}(t)-1\bigg\}.$ (6)
The task $z_{i}(t)$ will be immediately dropped if its processing is not completed by the end of the time slot $t+\Delta_{i}(t)-1$. In addition, we introduce $D_{i}^{\text{C}}(t)$ as the number of time slots required to complete the processing of task $z_{i}(t)$ on MD $i\in\mathcal{I}$.
@@
III-B2 Edge Execution
We model the edge execution by the queues associated with MDs deployed at ENs. If computation task $z_{i}(t^{\prime})$ is dispatched to EN $j$ in time $t^{\prime}<t$, we let $z_{i,j}^{\text{E}}(t)$, $\lambda_{i,j}^{\text{E}}(t)$ (in bits), and $\rho_{i,j}^{\text{E}}(t)$ denote the unique index of the task, the task size, and the number of CPU cycles required per unit of the task in the $i^{\text{th}}$ queue at EN $j$, respectively. We define $\eta_{i,j}^{\text{E}}(t)$ (in bits) as the length of this queue at the end of time slot $t\in\mathcal{T}$. We refer to a queue as an active queue in a certain time slot if it is not empty. That being said, if at least one task is already in the queue from previous time slots or there is a task arriving at the queue, that queue is active. We define $\mathcal{B}_{j}(t)$ to denote the set of active queues at EN $j$ in time slot $t$.
$\mathcal{B}_{j}(t)={\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\bigg\{i{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\big|\,i\in\mathcal{I},\lambda_{i,j}^{\text{E}}(t)>0\,\,{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\text{or}{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\,\,\eta_{i,j}^{\text{E}}(t-1)>0{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}\bigg\}.{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}i}$ (9)
@@ III-B2 Edge Execution
$\eta_{i,j}^{\text{E}}(t)\hskip-2.27621pt=\hskip-2.84526pt\left[\eta_{i,j}^{\text{E}}(t-1)\hskip-1.70717pt+\hskip-1.99168pt\lambda_{i,j}^{\text{E}}(t)\hskip-1.99168pt-\hskip-1.99168pt{f_{j}^{\text{E}}\tau\over\rho_{i,j}^{\text{E}}(t)b_{j}(t)}-\omega_{i,j}(t)\right]^{+}\hskip-2.84526pt.$ (10)
@@
$\hat{l}_{i,j}^{\text{E}}(t)=\max\bigg\{t,\max\limits_{t^{\prime}\in\{0,1,\ldots,t-1\}}l_{i,j}^{\text{E}}(t^{\prime})+1\bigg\},$ (11)
@@
$\sum_{t^{\prime}=\hat{l}_{i,j}^{\text{E}}(t)}^{l_{i,j}^{\text{E}}(t)}{f_{j}^{\text{E}}\tau\over\rho_{i,j}^{\text{E}}(t)b_{j}(t^{\prime})}\mathbbm{1}(i\in\mathcal{B}_{j}(t^{\prime}))\geq\lambda_{i,j}^{\text{E}}(t),$ (12)
$\sum_{t^{\prime}=\hat{l}_{i,j}^{\text{E}}(t)}^{l_{i,j}^{\text{E}}(t)-1}{f_{j}^{\text{E}}\tau\over\rho_{i,j}^{\text{E}}(t)b_{j}(t^{\prime})}\mathbbm{1}(i\in\mathcal{B}_{j}(t^{\prime}))<\lambda_{i,j}^{\text{E}}(t),$ (13)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D_{i,j}^{\text{E}}(t)=l_{i,j}^{\text{E}}(t)-\hat{l}_{i,j}^{\text{E}}(t).}$ (14)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{i,j}^{\text{E}}(t)=\sum_{t^{\prime}=\hat{l}_{i,j}^{\text{E}}(t)}^{l_{i,j}^{\text{E}}(t)}\hskip-5.69054pt{\,p_{j}^{\text{E}}\tau\over\,\,b_{j}(t^{\prime})\,\,}\mathbbm{1}(i\in\mathcal{B}_{j}(t^{\prime})),}$ (15)
where $p_{j}^{\text{E}}$ is constant and denotes the energy consumption of the EN $j$’s processor when operating at full capacity.
In addition to the energy consumed by EN $j$ for task processing, we also take into account the energy consumed by the MD $i$’s user interface in the standby state while waiting for task completion at the EN $j$. We define $E_{i}^{\text{I}}(t)$ as the energy consumption associated with the user interface of MD $i\in\mathcal{I}$, which is given by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E_{i}^{\text{I}}(t)=\sum_{j\in\mathcal{J}}y_{i,j}(t)D_{i,j}^{\text{E}}(t)p_{i}^{\text{I}}\tau,}$ (16)
where $p_{i}^{\text{I}}$ is the standby energy consumption of MD $i\in\mathcal{I}$. Recall that $y_{i,j}(t)$ is the binary offloading indicator. Moreover, among all $j\in\mathcal{J}$, $y_{i,j}=1$ only for one $j$ which is the corresponding EN.
IV Task Offloading problem Formulation
Based on the introduced system model, we present the computation task offloading problem in this section. Our primary goal is to enhance each MD’s QoE individually by taking the dynamic demands of MDs into account. To achieve this, we approach the optimization problem as an MDP, aiming to maximize the MD’s QoE by striking a balance among key QoE factors, including task completion, task delay, and energy consumption. To prioritize QoE factors, we utilize the MD’s energy level of the battery, which plays a crucial role in decision-making. Specifically, when an MD observes its state (e.g. task size, queue details, and battery status) and encounters a newly arrived task, it selects an appropriate action for that task. The selected action, based on the observed state, will result in enhanced QoE. Each MD strives to maximize its long-term QoE by optimizing the policy mapping from states to actions. In what follows, we first present the state space, action space, and QoE function, respectively. We then formulate the QoE maximization problem for each MD.
IV-A State Space
A state in our MDP represents a conceptual space that comprehensively describes the state of an MD facing the environment. We represent the MD $i$’s state in time slot $t$ as vector $\boldsymbol{s}_{i}(t)$ that includes the newly arrived task size, the queues information, the MD’s energy level of the battery, and the workload history at the ENs. The MD observes this vector at the beginning of each time slot. The vector $\boldsymbol{s}_{i}(t)$ is defined as follows:
@@ IV-A State Space
where vector $\boldsymbol{\eta}_{i}^{\text{E}}(t-1)=(\eta_{i,j}^{\text{E}}(t-1))_{j\in\mathcal{J}}$ represents the queues length of MD $i$ in ENs at the previous time slot and is computed by the MD according to (10). Since MDs are assumed to be battery-operated, their operation modes are aligned with those of real-world devices. Let $\phi_{i}(t)$ represent the battery level percentage of MD $i$ at time slot $t$, where $\phi_{i}(t)$ is selected from the discrete set $\Phi=\{\phi_{1},\phi_{2},\phi_{3}\}$, corresponding to ultra power-saving, power-saving, and performance modes, respectively.
In addition, to predict future EN workloads, we define the matrix $\mathcal{H}(t)$ as historical data, indicating the number of active queues for all ENs. This data is recorded over $T^{\text{s}}$ time slots, ranging from $t-T^{\text{s}}$ to t1t-1, in $T^{\text{s}}\times J$ matrix. For EN $j$ workload history at $i^{th}$ time slot from $T^{\text{s}}-t$, we define $h_{i,j}(t)$ as:
@@
IV-B Action Space
The action space represents the agent’s decisions. We define $\boldsymbol{a}_{i}(t)$ to denote the action taken by MD $i\in\mathcal{I}$ in time slot $t\in\mathcal{T}$. These actions involve two decisions, (a) Offloading decision to determine whether or not to offload the task, and (b) Offloading target to determine the EN to send the offloaded tasks. Thus, the action of MD $i$ in time slot $t$ can be concisely expressed as the following action tuple:
@@ IV-B Action Space
where vector $\boldsymbol{y}_{i}(t)=(y_{i,j}(t))_{j\in\mathcal{J}}$ represents the selected EN for offloading this task. In Section V-B, we will discuss the size of this action space.
IV-C QoE Function
The QoE function reflects user satisfaction with task computation, whether using local or edge resources. QoE accounts for task completion rate, processing delay, and energy consumption. Each user may have unique QoE requirements [40]. Therefore, we define a multi-dimensional, adaptive structure to assess QoE, balancing these factors according to the MDs preferences, such as prioritizing reduced task delay or energy saving, which may vary over time. To capture personalized QoE requirements, we define QoE as a weighted sum of the above factors. Each MD dynamically adjusts these weights to reflect the importance of different factors based on its energy modes, which are performance, power-saving, and ultra-power-saving modes. We now calculate task delay and energy consumption and then introduce the associated cost and QoE function.
Given the selected action $\boldsymbol{a}_{i}(t)$ in the observed state $\boldsymbol{s}_{i}(t)$, we represent $\mathcal{D}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$ as the delay of task $z_{i}(t),$ which indicates the number of time slots from time slot $t$ to the time slot in which task $z_{i}(t)$ is processed. It is calculated by:
$\mathcal{D}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))=(1-x_{i}(t))\Big(l_{i}^{\text{C}}(t)-t+1\Big)$
$\hskip 8.5359pt+\,x_{i}(t)\Bigg(\sum\limits_{j\in\mathcal{J}}\sum\limits_{t^{\prime}=t}^{T}\mathbbm{1}\big(z_{i,j}^{\text{E}}(t^{\prime})=z_{i}(t)\big)\,l_{i,j}^{\text{E}}(t^{\prime})-t+1\Bigg),$ (20)
where $\mathcal{D}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))=0$ if task $z_{i}(t)$ is dropped. Correspondingly, we denote the energy consumption of task $z_{i}(t)$ when taking action $\boldsymbol{a}_{i}(t)$ in the observed state $\boldsymbol{s}_{i}(t)$ as $\mathcal{E}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$, which is
$\mathcal{E}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))=(1-x_{i}(t))E_{i}^{\text{L}}(t)+x_{i}(t)\bigg(E_{i}^{\text{T}}(t)+$
$\hskip 34.14322ptE_{i}^{\text{I}}(t)+\sum\limits_{j\in\mathcal{J}}\sum\limits_{t^{\prime}=t}^{T}\mathbbm{1}\big(z_{i,j}^{\text{E}}(t^{\prime})=z_{i}(t)\big)E_{i,j}^{\text{E}}(t^{\prime})\bigg).$ (21)
To define associated cost, we use a weighted sum of task delay $\mathcal{D}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$ and energy consumption $\mathcal{E}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$, where the MD dynamically adjusts the weights based on its energy level to reflect the preference for each factor. Given the delay and energy consumtion of task $z_{i}(t)$, we also define $\mathcal{C}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$ that denotes the assosiate cost of task $z_{i}(t)$ given the action $\boldsymbol{a}_{i}(t)$ in the state $\boldsymbol{s}_{i}(t)$.
@@ IV-C QoE Function
where $\phi_{i}(t)$ represents the MD $i$’s energy level. When the MD is operating in performance mode, the primary focus is on minimizing task delays, thus the delay contributes more to the cost. On the other hand, when the MD switches to ultra power-saving mode, the main attention is directed toward reducing power consumption.
Finally, we define $\boldsymbol{q}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))$ as the QoE associated with task $z_{i}(t)$ given the selected action $\boldsymbol{a}_{i}(t)$ and the observed state $\boldsymbol{s}_{i}(t)$. The QoE function is defined as follows $\boldsymbol{q}_{i}(\boldsymbol{s}_{i}(t),\boldsymbol{a}_{i}(t))=$
[…]